OpenStudy (anonymous):

This equation, x2 + y2 - 6x = 0, represents:

myininaya (myininaya):

a circle

myininaya (myininaya):

x^2-6x+y^2=0 x^2-6x+9+(y-0)^2=9 (x-3)^2+(y-0)^2=9 circle with center (3,0) and radius 3